OpenStudy (anonymous):

The sum of two numbers is 36. The sum of the squares of the two numbers is 720. What is the product of the two numbers?

OpenStudy (anonymous):

x+y = 36 x^2 = y^2 = 720 solve, (12,24) or (24,12) 12*24 = 288

OpenStudy (anonymous):

i meant x^2 + y^2 = 720

OpenStudy (anonymous):

thank you